9.65 C of electric current is passed through fused anhydrous magnesium chloride. The magnesium metal thus, obtained is completely converted into a Grignard reagent. The number of moles of the Grignard reagent obtained is

Grade:12

1 Answers

Arun
6 years ago
W = Z*I*t
Where Z = eq. Weight /F = 12/96500
I *t = 9.65C
W = 12* 9.65/96500 = 12 * 10^-4 gm
 
24 gm of Mg+2 = 1 mol of Mg+2
Hence
12* 10^-4 gm of Mg+2 = 12* 10^-4/ 24
 = 5* 10^-5 moles